### v1

# NoveltyRank: Estimating Conceptual Novelty of AI Papers

NoveltyRank:
Estimating Conceptual Novelty of AI Papers
Zhengxu Yan1,*, Han Li2,*, and Yuming Feng2,*
1Department of Computer Science, Stanford University
2Department of Electrical Engineering, Stanford University
jasonyan@stanford.edu, cli27@stanford.edu, yumingf@stanford.edu
1 Introduction
With the growing ease of academic publishing, the volume of research papers, especially in AI-related fields, has surged dramatically. This "flood" of publications make it difficult for truly novel and impactful work to stand out, and manual novelty assessment is often unstable and time-consuming. Our project aims to develop a model that estimates and ranks the conceptual novelty of AI papers, enabling a data-driven and scalable assessment of research originality. Such a system can help researchers efficiently identifying submissions that introduce genuinely innovative ideas rather than minor variants, and provide conference reviewers with a quantitative and consistent signal of novelty.
Our approach evaluates novelty primarily through a paper’s title, abstract, and semantic similarity to prior literature. Given the motivation of novelty estimation, we explore two task formulations with different modeling objectives, each offering a different perspective: (1) Binary classifcation, which predicts the paper’s absolute novelty from learned patterns of prior novel works, and (2) Pairwise novelty comparison, which learns to distinguish papers by relative novelty over others. We fine-tune Qwen3-4B-Instruct-2507 qwen and SciBERTscibert on both tasks, benchmarking against GPT-5.1 to analyze how task formulation and modeling choices affect performance. The implementation is publicly available at https://github.com/ZhengxuYan/NoveltyRank.
2 Dataset and Features
2.1 Description
Our dataset is built from web-scraped data and the public ICLR 2017–2025 dataset gonzalez2024learning, totaling 60,294 papers published between 2023 and 2025: 50,442 randomly sampled arXiv papers and 9,852 accepted by top-tier venues across six domains (AI, ML, CV, Robotics, NLP, and Cryptography). Each entry includes metadata such as paper ID, publication date, authors, title, and abstract.
Conference acceptance serves as a proxy for novelty, assuming accepted papers generally reflect more original or impactful contributions: Accepted papers are labeled as positives (1), while randomly sampled arXiv papers are negatives (0). To prevent temporal leakage, we split data by publication date: Training on papers from 2024 to early 2025, and testing on papers after March 15, 2025. This setup ensures evaluation on future, unseen publications, aligning with real-world scenarios.
2.2 Embedding
We embed each paper’s title and abstract using SPECTER2 specter2, a transformer fine-tuned for scientific document representation. We employed two adapter modules:
  1. 1.
    Classification Embedding – captures a paper’s semantic identity, such as topic, technical depth, and conceptual contribution.
  2. 2.
    Proximity Embedding – trained with citation-based contrastive learning to measure closeness in citation space. Based on proximity embedding, we adopt Faiss faiss for scalable nearest-neighbor search and compute papers’ similarity scores given their top-10 similar prior literature (rolling by publication date), serving as additional engineered features.
Combining these embeddings enables the model to assess both content-level representation and contextual similarity for novelty evaluation.
3 Models
To implement the NoveltyRank framework, we formulate novelty estimation through two distinct task both aligned with the goal of evaluating scientific innovation: Binary Classification (evaluating the absolute novelty of a single paper) and Pairwise Comparison (assessing relative novelty between two papers). These tasks differ in input structure and supervision.
To benchmark performance across computational scales, we experiment each task formulation on three models listed below following a specific logic: as model size decreases, the degree of task-specific adaptation increases. Such comparison determine whether smaller models with targeted parameter updates can match or surpass larger general-purpose models, verifying the feasibility of lightweight deployment.
  • GPT-5.1 (Large-Scale / Zero-shot): As a frontier model, GPT-5.1 serves as the upper-bound baseline. It is accessed via API and evaluated in zero-shot without parameter updates.
  • Qwen3-4B (Mid-Scale / LoRA Tuning): Qwen3-4B balances size and adaptability. We apply a two-stage fine-tuning with Supervised Fine-Tuning (SFT) followed by Direct Preference Optimization (DPO) dpo, using LoRA lora to update a small subset of parameters while freezing the backbone.
  • SciBERT (Small-Scale / Layer-Frozen Fine-Tuning): SciBERT represents the compact, domain-specific model. We adopt a multimodal approach by concatenating the standard SciBERT [CLS] token with pre-computed SPECTER2 embeddings and similarity features. To preserve scientific linguistic knowledge, the lower 8 encoder layers are frozen, and only upper layers and task-specific heads are fine-tuned.
4 Task Formulation 1: Binary Novelty Classification
For binary novelty classification, we formulate the task as a supervised learning problem where the model primarily use papers’ title, abstract, and similarity scores to predict binary novelty label (0 or 1). The objective is to learn general patterns and absolute criteria from established novel papers to assess the novelty of unseen samples.
4.1 Qwen3-4B: SFT and DPO Fine-tuning
Figure 1: Qwen3-4B Pipeline for Binary Classification
The full pipeline of Qwen3-4B model for binary novelty classification task is illustrated in Fig. 1. Because novelty is inherently comparative, we augment this input with a Similarity Report generated by an LLM Judge, which summarizes overlaps and distinctive contributions relative to the top-$K$ similar papers retrieved via FAISS faiss. This comparison serves provides an additional signal for decision-making process.
Although a base Qwen3-4B model could serve as the judge, we use Qwen-235B to exploit its superior reasoning capabilities for higher-quality analysis. To optimize training efficiency, all similarity reports are pre-generated. We further improve judge performance through prompt engineering, including Chain-of-Thought (CoT) chain and few-shot examples shot. A prompt example is provided in the Appendix.
Supervised Fine-Tuning (SFT)
In SFT, we apply cross-entropy loss between the model’s generated tokens and ground-truth labels. This setup naturally yields a binary supervision signal, training the model to generate discrete outputs (0 or 1). We avoid prompting for continuous confidence scores (e.g., 0.85), as LLMs generate text tokens rather than mathematically grounded probabilities. The decimals ouputs would be linguistic hallucinations, rendering the model unreliable for quantitative confidence estimation.
Direct Preference Optimization (DPO)
For DPO, we initialize the model using the SFT checkpoint to ensure training stability. We construct preference pairs $(y_{chosen},y_{rejected})$ based on the ground truth of submission acceptance: if the chosen response is "1" and the rejected response is "0", the correct label is "1" (and vice versa). DPO then optimizes the model’s likelihood to favor the correct classification over the incorrect one, further refining the model’s ability to robustly distinguish novelty.
4.2 Fine-tuned SciBert
We fine-tune the pretrained scibert-scivocab-uncased model for binary classification by integrating textual and metadata features. The input sequence combines the paper’s title, abstract, and primary categories, separated by [SEP] tokens and truncated to a maximum length of 512. To capture semantic context beyond the input text, we concatenate the SciBERT [CLS] token output (768-dim) with three external feature vectors: the SPECTER2 classification embedding (768-dim), the proximity embedding (768-dim), and an aggregated embedding of the top-10 similar papers (768-dim).
These features are fused into a multi-modal representation and passed through a custom classification head consisting of a three-layer feed-forward network ($2306\to 512\to 128\to 2$) with ReLU activation and a dropout rate of 0.1. The model is trained using Cross-Entropy loss with an AdamW optimizer ($\eta=2e^{-5}$) and a linear warmup scheduler for 5 epochs.
4.3 Evaluation Metrics
Given binary classification task and the class imbalance in our dataset, we evaluate model performance primarily using Precision, Recall, and F1-Score, which better reflect effectiveness on the minority (novel) class. Accuracy is also reported for completeness, but it may be misleading in cases where the model predicts predominantly negative labels.
4.4 Results and Discussion
Table 1 presents the performance of binary classification task on the test set, which consists of 10,889 examples with a highly imbalanced distribution (1,358 positives, approximately 12.5%). The discussion follows.
Table 1: Test Performance of Binary Classification
Model Accuracy Precision Recall F1-score
GPT-5.1 0.242 0.120 0.986 0.215
SFT Qwen3-4B 0.627 0.194 0.632 0.297
DPO Qwen3-4B 0.612 0.205 0.735 0.321
Fine-tuned SciBert 0.744 0.187 0.313 0.234
Performance of Large-Scale Models
GPT-5.1 exhibits a strong "generosity bias" (Recall 0.986, Precision 0.120). Without specific training, the model tends to label nearly all papers as novel, failing to establish a rigorous boundary to filter out incremental works.
Effectiveness of DPO
DPO improves upon the SFT baseline (F1 0.321 vs. 0.297) primarily by boosting Recall (0.735 vs. 0.632). This confirms that preference optimization effectively encourages the model to identify valid novelty signals actively rather than defaulting to safe predictions.
The Accuracy Paradox
Despite higher accuracy (e.g., SciBERT’s 74.4%), fine-tuned models suffer from dataset imbalance. Low recall (especially SciBERT’s 0.313) indicates they models collapse into conservative classifiers, minimizing loss by over-predicting the majority "non-novel" class.
Insights on Absolute Novelty
The suboptimal F1-scores across all models suggest that novelty is inherently relative, not absolute. Learning a crisp binary boundary from isolated inputs is difficult due to vague definitions. This limitation motivates our shift to a pairwise comparison formulation, where novelty is assessed relatively rather than absolutely.
Figure 2: Siamese SciBERT Network
5 Task Formulation 2: Pairwise Novelty Comparison
For pairwise novelty comparison, the model jointly evaluates two papers to determine which one exhibits greater originality. Unlike binary classification, which learns absolute novelty criteria, this formulation focuses on relative patterns of innovation, enabling the model to discriminate novelty based on direct comparison.
Although each paper uses the same feature set as in the binary classification task, we construct comparison pairs to reflect the comparative nature of the objective. For each novel paper, we sample a non-novel counterpart from the same domain to ensure meaningful contrast. To address class imbalance, we generate five such comparison pairs per positive paper by randomly sampling multiple negative examples. To prevent positional bias (e.g., model systematically favoring the first option), we randomly shuffle the order of two papers within the pair during training.
5.1 Qwen3-4B: SFT and DPO Fine-tuning
The same prompt-engineering methods and the SFT and DPO methods described in Section 4.1 applies here. However, the supervision signal shifts from binary labels (0 or 1) to positional indicators (Paper A or Paper B). This allows the model to adapt the same optimization pipeline to a relative comparison setting.
5.2 Siamese SciBERT Network
To support the pairwise comparison task, we implement a Siamese network architecture with shared weights (as shown in Fig. 2). The model takes a pair of papers $(P_{A},P_{B})$ as input, processing each through identical SciBERT encoders to produce scalar novelty scores $s_{A}$ and $s_{B}$. To improve generalization and prevent overfitting on the smaller dataset of pairs, we freeze the embeddings and the first 8 transformer layers, fine-tuning only the top 4 layers of the encoder.
Similar to the classification task, the textual representation is concatenated with classification and proximity embeddings. This combined vector is fed into a scoring head ($2306\to 256\to 64\to 1$) with a higher dropout rate of 0.5 to act as a regularizer. The network is optimized using RankNet loss, formulated as binary cross-entropy on the score difference $\sigma(s_{A}-s_{B})$, effectively maximizing the likelihood that the novel paper is scored higher than its non-novel counterpart.
5.3 Evaluation Metrics
For the novelty comparison task, we evaluate performance using Pairwise Agreement, the proportion of pairs in which the model correctly identifies the more novel paper. Unlike the training phase which relied on random sampling (1:5 ratio), evaluation employs a dense pairing strategy: every positive paper is paired with all available negative samples in the same domain. This exhaustive matching eliminates sampling variance and maximizes the utilization of the test set for a robust assessment.
5.4 Results and Discussion
We constructed 9,531 testing pairs spanning six distinct domains. Table 2 presents the aggregate test agreement rates, while Figure 3 provides a detailed breakdown of the test performance by domain and illustrates its distribution within the training set.
Table 2: Performance of Pairwise Comparison
Metric GPT-5.1 SFT Qwen3-4B DPO Qwen3-4B FT SciBERT
Agreement 0.583 0.739 0.741 0.753
Efficacy of Fine-Tuning and Task Formulation
The results demonstrate that task-specific fine-tuning offers a clear advantage over generalized large-scale models. While the GPT-5.1 baseline achieved only marginal agreement (0.583), all fine-tuned models performed substantially better, led by SciBERT (0.753) and DPO-tuned Qwen3-4B (0.741).
These significant gains validate two key points: First, that domain-adapted models outperform frontier models for our task, despite their smaller size. Second, the high, consistent agreement rates confirm the effectiveness of the pairwise comparison formulation, which provides a clearer, more actionable training signal than binary classification formulation.
Figure 3: Comparison performance by domain. The bar chart shows each domain’s test agreement rates alongside the proportion of positive labels and the category distribution within the training set.
The "Consistency over Quantity" Paradox
A closer examination of category-specific performance (Figure 3) reveals an inverse relationship between data volume and model accuracy. Models achieve their highest agreement in small, imbalanced fields such as Robotics (0.81–0.84) and Cryptography, even though these categories constitute only a minor portion of the training data. In contrast, large fields like Machine Learning yield lower agreement.
This suggests that pair consistency, rather than dataset size, is the primary driver of optimization quality. Niche domains tend to be semantically compact, producing pairs where “novel” and “non-novel” examples are closely aligned and easier to compare. Large heterogeneous domains, however, generate pairs spanning divergent subtopics, reducing semantic coherence and making relative novelty harder to judge. Overall, these results indicate that for pairwise comparison tasks, well-structured, semantically aligned pairs matter more than raw data scale.
6 Conclusion
This project introduces NoveltyRank, a system to evaluatw the conceptual originality of AI papers. We find that pairwise comparison formulation offers a significantly cleaner and more effective learning signal than binary classification formulation. Moreover, We investigate three model scales and find that the cost-effective, domain-specific SciBERT, fine-tuned under the comparison setting, achieves best results. This demonstrates that comparative structure with small-model fine-tuning provides a more efficient and effective solution than increasing model size.
Contributions
All team members contributed to project deliverables.
Jason Yan: Conceived the initial project idea and developed the data scraping pipeline. Implemented the Siamese SciBERT multimodal network, and engineered the user interface for demonstration.
Christine Li: Generated SciBERT and Specter2 text embeddings and performed the FAISS similarity search. Conducted the downstream similarity computation and performed data validation tasks.
Yuming Feng: Architected the core comparison task formulation and implemented the entire end-to-end training pipeline for Qwen3-4B using both SFT and DPO methodologies. Primarily authored the Interpretation and Discussion of the experimental results.
Acknowledgments
We would like to express our sincere gratitude to Cody Ho and Ryan Lian for their helpful feedback and guidance during the course of this project.
Appendix
A. Deployment
The best-performed model is deployed in a user-friendly web application [^1] https://novelty-rank.vercel.app/ that allows interactive exploration of novelty evaluation and ranking.
The complete dataset created for this project is available on HuggingFace [^2] https://huggingface.co/datasets/JasonYan777/novelty-rank-with-similarities to support reproducibility and future research.
B. Hyperparameters
Key hyperparameters for Qwen3-4B (SFT and DPO) and SciBERT includes:
  • Number of epochs, batch size, and learning rate
  • Optimization settings
  • LoRA rank and scaling parameters (for Qwen3-4B)
B.1 Qwen-4B Hyperparameters
Table 3: Hyperparameters for Qwen3-4B across classification and comparison tasks (SFT and DPO).
Hyperparameter Class-SFT Class-DPO Comp-SFT Comp-DPO
Learning rate 2e-5 1e-6 3e-5 1.5e-6
Batch size 256 128 64 64
Epochs 4 1 10 4
Max sequence length 4096 1024 4096 1024
LR scheduler linear linear linear linear
Optimizer AdamW AdamW AdamW AdamW
Adam $\beta_{1}$ 0.9 0.9 0.9 0.9
Adam $\beta_{2}$ 0.95 0.95 0.95 0.95
Adam $\epsilon$ 1e-8 1e-8 1e-8 1e-8
LoRA rank ($r$) 32 32 32 32
DPO $\beta$ (only for DPO) 0.1 0.1
B.2 SciBERT Hyperparameters
Table 4: Hyperparameters for SciBERT across binary classification and pairwise comparison tasks.
Hyperparameter Binary Class. Pairwise Comp.
Learning rate 2e-5 1e-5
Batch size 32 64
Epochs 5 5
Max sequence length 512 512
Weight decay 0.01 0.1
Warmup ratio 0.1 0.1
Optimizer AdamW AdamW
Gradient accumulation 1 2
Dropout rate 0.1 0.5
Frozen layers None Embeddings + Layers 0-7
C. Prompt Templates
C.1 Binary Novelty Classification Prompt
# System Prompt
You are an expert AI researcher and senior conference reviewer (NeurIPS/ICLR level).
Your goal is to judge whether the submission introduces a conceptually novel idea.
Conceptual novelty captures fundamental shifts in scientific thinking.
### Conceptual Novelty Primer
Consider the following signals:
- Problem Formulation: Does it redefine an existing task or introduce a new one?
- Methodological Innovation: Does it propose a new class of algorithms or training paradigm?
- Theoretical Insight: Does it deliver a unifying or surprising theoretical lens?
- Cross-Disciplinary Import: Does it import a transformative idea from another domain?
Incremental tweaks (hyperparameters, surface-level architecture edits, dataset swaps) are not novel.
### Reference Decisions
Example 1:
Title: Differentiable Logic for Robotics
Abstract: Introduces a framework that composes continuous control policies with symbolic logic programs to enable reasoning-guided motion planning.
Similarity scores: max=0.61 | avg=0.48
Reasoning: Combines two previously disjoint paradigms (continuous control and symbolic reasoning) into a unified differentiable architecture (Novel).
Output: 1
Example 2:
Title: Better Hyperparameters for BERT Fine-Tuning
Abstract: Reports extensive sweeps over learning rates and batch sizes for BERT on GLUE benchmarks.
Similarity scores: max=0.89 | avg=0.81
Reasoning: Purely empirical tuning without a new formulation or architecture (Not Novel).
Output: 0
Example 3:
Title: Physical Priors for Diffusion Models
Abstract: Incorporates symbolic conservation laws into diffusion model training to improve controllable generation.
Similarity scores: max=0.67 | avg=0.58
Reasoning: Introduces a cross-disciplinary inductive bias that reshapes the generative objective (Novel).
Output: 1
# User Prompt
### Paper Metadata
Title: {title}
Primary Category: {category}
Abstract: {abstract}
Max similarity to prior work: {max_sim}
Average similarity to prior work: {avg_sim}
### Similarity Report (Aggregated)
{similarity_report}
### Decision Instructions
1. Synthesize the available evidence (abstract + similarity signals).
2. Decide whether the work represents a conceptually novel contribution.
3. Output "1" if the paper is conceptually novel and likely to influence future research.
4. Output "0" if the contribution is incremental, derivative, or lacks conceptual novelty.
Respond with a single digit (0 or 1).
C.2 Pairwise Novelty Judgment Prompt
# System Prompt
You are an expert computer-vision researcher and senior conference reviewer (CVPR/ICCV/NeurIPS level).
Your goal is to compare the *conceptual novelty* of two computer-vision research papers (not just surface/benchmark improvements).
Conceptual Novelty Primer
Consider the following signals:
- Problem Formulation: Does it redefine an existing task or introduce a new one?
- Methodological Innovation: Does it propose a new class of algorithms or training paradigm?
- Theoretical Insight: Does it deliver a unifying or surprising theoretical lens?
- Cross-Disciplinary Import: Does it import a transformative idea from another domain?
Incremental tweaks (hyperparameters, surface-level architecture edits, dataset swaps) are not novel.
Step-by-step reasoning (use these as your guide and mention the strongest signal):
1) Extract the core technical idea from each papers title and abstract.
2) Check whether the idea represents a new task, representation, learning paradigm, or major architectural shift.
3) Use similarity metrics as supportive evidence (high similarity tilts toward incremental), but prioritize conceptual signals (new objective, representation, or theory).
4) Choose which paper is more conceptually novel; answer only with A or B’.
--- EXAMPLES
Example 1:
Paper A: Introduces Vision Transformer (ViT), treats images as a sequence of patches and applies a pure transformer backbone, changing core architecture for vision.
Paper B: Reports small regularization and augmentation tweaks to ResNet training that marginally improve accuracy.
Reasoning: A introduces a new architectural paradigm for visual representation (Novel).
Output: A
Example 2:
Paper A: Proposes Neural Radiance Fields (NeRF), an implicit continuous 3D scene representation enabling view synthesis.
Paper B: Improves an existing multi-view stereo pipeline with a better post-processing filter.
Reasoning: NeRF introduces a fundamentally new representation and rendering paradigm (Novel).
Output: A
Example 3:
Paper A: Applies an off-the-shelf transformer to a small medical imaging dataset with minor changes.
Paper B: Proposes a new contrastive objective that aligns multi-resolution feature maps and demonstrates broad transfer across many vision tasks.
Reasoning: B defines a new learning objective with broad implications -> Novel.
Output: B
# User Prompt
### Paper A
Title: {titleA}
Primary Category: {categoryA}
Abstract: {abstractA}
Max similarity to prior work: {max_simA:.4f}
Average similarity to prior work: {avg_simA:.4f}
### Paper B
Title: {titleB}
Primary Category: {categoryB}
Abstract: {abstractB}
Max similarity to prior work: {max_simB:.4f}
Average similarity to prior work: {avg_simB:.4f}
Output only the single letter A or B’.